h(x)=1/x,find h(- 1/2)
Answer by tutorcecilia(2152)

h(x)=1/x,find h(- 1/2)
Plug-in x=-1/2 and solve
h(-1/2)=1/(-1/2)
=(1)(-2/1)=-2